# PEAR: Position-Embedding-Agnostic Attention Re-weighting Enhances Retrieval-Augmented Generation with Zero Inference Overhead

### v1

@@ Abstract
Large language models (LLMs) enhanced with retrieval-augmented generation (RAG) have introduced a new paradigm for web search. However, the limited context awareness of LLMs degrades their performance on RAG tasks. Existing methods to enhance context awareness are often inefficient, incurring time or memory overhead during inference, and many are tailored to specific position embeddings. In this paper, we propose Position-Embedding-Agnostic attention Re-weighting (PEAR), which enhances the context awareness of LLMs with zero inference overhead. Specifically, on a proxy task focused on context copying, we first detect heads which suppress the models’ context awareness, thereby diminishing RAG performance. To weaken the impact of these heads, we re-weight their outputs with learnable coefficients. The LLM (with frozen parameters) is optimized by adjusting these coefficients to minimize loss on the proxy task. As a result, the coefficients are optimized to values less than one, thereby reducing their tendency to suppress RAG performance. During inference, the optimized coefficients are fixed to re-weight these heads, regardless of the specific task at hand. Our proposed PEAR offers two major advantages over previous approaches: (1) It introduces zero additional inference overhead in terms of memory usage or inference time, while outperforming competitive baselines in accuracy and efficiency across various RAG tasks. (2) It is independent of position embedding algorithms, ensuring broader applicability.
@@ 1 Introduction
Retrieval-augmented generation (RAG, (Lewis et al., 2021)) is widely utilized to enhance large language models (LLMs) on tasks like question answering. Typically, an RAG framework retrieves documents related to users’ question from external knowledge bases or web pages, and then arranges them in the LLMs’ context as the references to form answers. This LLM-based question-answering paradigm has given rise to a promising web search paradigm (Microsoft, 2023; OpenAI, 2024).
Recent research demonstrated LLMs’ limitations on context awareness, especially when processing long context. These limitations in LLMs’ context awareness challenge the effectiveness and robustness of RAG frameworks. For instance, Liu et al. (2023) found that when performing in-context retrieval tasks, LLMs exhibit insensitivity to information located in the middle of the context, a phenomenon referred to as “lost in the middle.” Chen et al. (2024) identified a mathematical property in rotary position embedding (RoPE, (Su et al., 2023)) that results in LLMs assigning less attention to specific contextual positions, leading to varying context awareness throughout the entire context.
Existing approaches to enhancing LLMs’ context awareness are inefficient in terms of memory and time cost. Some works (Peysakhovich & Lerer, 2023) segment and re-arrange the input context, with the assumption that placing important information in positions the model attends well can improve RAG’s effectiveness. This method incur additional inference time costs, negatively affecting user experience, as it requires multiple forward passes to obtain attention weights for guiding segment rearrangement. Another group of studies modifies the model’s working mechanism, specifically by employing a set of position embeddings to adjust the attention preferences of attention heads. While these methods are input-agnostic, they also lack efficiency due to “parallelable” forward passes (Chen et al., 2024) although without increased time cost, disrupting the parallelism of multi-head attention (resulting in increased time cost) to achieve low memory cost (Lin et al., 2024), or still requiring “non-parallelable” multiple forward passes (Zhang et al., 2024) which is similar to (Peysakhovich & Lerer, 2023). Moreover, these studies are mainly designed for RoPE and face challenges in generalizing to other position embedding algorithms, limiting their broader applications.
@@
    Prior research (McDougall et al., 2023; Lv et al., 2024a) detected some attention heads decrease the language model’s prediction confidence by suppressing the flow of contextual information to the final position within the context, where the output is to be generated.
@@
As a result, we contend that such suppression mechanism in LLMs can be safely111“Safely” means that the parametric knowledge and fundamental capabilities remain unaffected. Detailed experimental results are presented in Section 5.4. weakened in RAG scenarios, thereby improving the performance of RAG. Our proposed PEAR includes two stages:
@@
In the second stage, we weaken detected RAG-suppression heads by re-weighting their outputs using learnable coefficients. These coefficients are optimized by minimizing the LLM’s loss on the proxy task, with the objective of next-token prediction in a supervised fine-tuning process (loss is computed only for the second half of the random-token sequence). During the optimization, the original LLM parameters are frozen. Intuitively, the learned coefficients are optimized to values less than one, reducing the relative weight of these heads compared to others in the same layer when multi-head outputs are aggregated. Consequently, their influence during the forward pass are weakened. Once optimized, the coefficients remain fixed and are agnostic to downstream RAG tasks.
PEAR achieves three-fold contributions:
@@ 2.1 Context awareness enhancement
Many studies highlighted limitations in LLMs’ context awareness. For example, Lu et al. (2022) found that the order of in-context learning (ICL) demonstrations significantly affects ICL accuracy. Liu et al. (2023) demonstrated that LLMs exhibit stronger awareness of content at the beginning and end of context but weaker awareness in the middle, a phenomenon termed “lost in the middle.” Chen et al. (2024) proposed that LLMs’ context awareness fluctuates across token positions due to mathematical properties in position embeddings. These challenges impact applications like RAG that rely on robust context awareness.
@@
Our proposed method, PEAR, enhances context awareness by weakening the RAG-suppression heads during the forward pass. It introduces no additional modules or extra forward passes, resulting in zero additional overhead in memory usage and inference time. Additionally, PEAR operates independently of position embedding algorithms, which is applicable to more LLMs compared to existing approaches.
@@ 2.2 Mechanistic interpretability
Investigating the role of a specific head during a forward pass is one of key focuses in mechanistic interpretability research. Wang et al. (2023) reported that in GPT-2 small (Radford et al., 2019), the 7th head in the 10th layer, termed the “negative head,” significantly hinders answer copying from context. McDougall et al. (2023) comprehensively studied this head, suggesting it functions as a self-repair mechanism to prevent overconfident outputs. Lv et al. (2024a) found that negative heads exist across various LLMs, employing different mechanisms to mitigate overconfidence, such as generating counteracting vectors or introducing high-frequency tokens’ information. This paper does not examine what specific mechanisms the heads employ to suppress RAG performance but instead aims to discover and suppress heads negatively impactful across general RAG tasks.
Yu et al. (2023) detected two types of heads in Transformer-based language models during counter-factual task execution (where counter-factual knowledge is provided in the context): memory heads, which prefer to use stored knowledge, and in-context heads, which prefer to use facts in the context. However, when re-weighting these heads, only reducing the weight of memory heads successfully enhances the model’s preference of using contextual knowledge, while enhancing or mitigating in-context heads does not bring much influence. Moreover, existing results only demonstrate effectiveness of these heads in the “country-capital” task (i.e., prompting the model to answer the capital city of the given country). Additionally, there is no evidence suggesting that re-weighting these heads improves the comprehensive context awareness of LLMs. We owe these failures to their head detection method, and the same re-weighting value applied for all the heads of the same type. By contrast, in our proposed PEAR, each individual head is re-weighted by a specific learnable coefficient, which is optimized through a proxy task independent of downstream tasks.
@@ 3 Preliminaries: Discovery of influential attention heads
    We conduct an intervention on a particular head $A^{(l,h)}$ at a specific position $i$ (e.g., the country token position in above examples) in the normal run by substituting its outputs with $\tilde{\textbf{a}}^{(l,h)}_{i}$. The subsequent activations in the computational graph are then recomputed (these reccomputed activations are denoted as $\bar{\textbf{a}}$ in the figure). If the final output of the language model as the intervention expects (e.g., the predicted token changes from ”Paris” to ”London”), the head $A^{(l,h)}$ is considered to have a positive influence on the processing of sequence $X$.
@@ 4.1 Discovery of RAG-suppression heads
For each input sample, we create a sequence of length $n$, denoted as $\{x_{1},\ldots,x_{n}\}$, where each $x_{i}$ is a randomly sampled token from the vocabulary. This sequence is repeated to form an input sample $X=\{x_{1},\ldots,x_{2n}\}$, with $x_{i}=x_{i+n}$ for $i\in[1,n]$. Research has shown that, in semantically meaningless contexts, models tend to check if the last few tokens in the sequence appeared previously and copy the suffix of their last appearance as the output (Olsson et al., 2022; Lv et al., 2024b). We consider an arbitrary LLM to successfully perform the proxy task when, at position $n+i+1$, the token with the highest output logits is $x_{i}$. Table 1 shows an example input.
@@
    The random token composition in $X$ ensures semantically meaningless input, minimizing knowledge bias and thereby enabling the discovered attention heads to have general RAG-related functions, independent of specific downstream tasks.
@@
    We focus on detecting changes in logits at position $2n$, where the model is expected to copy the token from position $n-1$. Consequently, we intervene at $a^{(l,h)}_{n-1}$ by replacing it with the saved mean vectors.
@@
    $\Delta\pi^{(l,h)}=\frac{\tilde{\pi}^{(l,h)}_{2n}[x_{n-1}]}{\pi_{2n}[x_{n-1}]}-1,$ (1)
    where $\pi_{2n}$ represents the final logits at position $2n$ during the normal run, and $[x_{n-1}]$ denotes selecting the value of the token $x_{n-1}$ from the logits. $\tilde{\pi}^{(l,h)}$ indicates the logits after intervention on $A^{(l,h)}$. We contend that a higher value of this metric suggests a stronger suppression effect from $A^{(l,h)}$.
@@
Notably, we do not suggest that the heads we discovered suppressing RAG tasks operate through the same mechanisms. These heads may serve various functions, such as copy suppression (McDougall et al., 2023), incorporating high-frequency token information (Lv et al., 2024a), or influencing the behavior of other heads to indirectly affect outputs (Wang et al., 2023). Analyzing the specific mechanism for each head is not the focus of this paper and is left for future works.
@@ 4.2 Re-weighting coefficient learning
In theory, our approach, PEAR, introduces zero additional overhead during inference on downstream RAG tasks, as it does not incorporate extra computational modules; instead, it only adjusts the aggregation weights of specific heads. In practice, however, the re-weighting process involves an additional multiplication, which results in zero additional overhead in inference time and memory usage. Additionally, the learning of re-weighting coefficients is independent of the LLM architecture, thus making our method compatible with various position embedding algorithms.
@@ 5.3 Applicability to LLMs using various position embeddings
In this section, we demonstrate the applicability of PEAR to LLMs utilizing different position embeddings. We conduct a multi-document question-answering (MDQA) experiment based on data from (Liu et al., 2023). Following Liu et al. (2023); Chen et al. (2024), we position the gold document (i.e., the document contains the ground truth answer) at various contextual positions to evaluate the robustness of a context-awareness enhancement method. In our experiments, we set the maximum document count to 10 and assess the question-answering accuracy when the gold document is placed as the 1st, 3rd, 5th, 7th, and 10th document, respectively. Since baseline methods are not compatible with the OPT and Baichuan models, we compare PEAR only with the original models. The results are presented in Table 5.
@@ 5.4 PEAR does not diminish knowledge capabilities in LLMs
Previous research (Geva et al., 2023; Lv et al., 2024a) has shown that certain attention heads store or play a crucial role in eliciting parametric knowledge. This raises the question of whether PEAR enhances context awareness of LLMs at the expense of their ability to utilize this parametric knowledge. To investigate this, we evaluated a PEAR-enhanced Llama2-7B-chat model using the MMLU benchmark (Hendrycks et al., 2021), and the results are presented in Table 6. The performance of the enhanced Llama2-7B-chat and the original Llama2-7B-chat did not show a significant difference. Consequently, we argue that PEAR, through its precise head discovery and effective re-weighting learning approaches, does not compromise the knowledge capabilities of LLMs.
@@
5.5 Analysis: The effect of $K$
While we have demonstrated the effectiveness of PEAR from various angles, a key point for discussion is the role of $K$, representing the number of heads to re-weight. Using Llama2-7B-chat as a case study, we vary $K$ and observe its impact on PEAR’s performance. Table 7 presents results on RAG tasks, while Table 8 details analysis for MDQA tasks. The findings indicate that PEAR performs optimally when $K$ matches the inherent threshold of the model, i.e., the number of heads with a significantly higher $\Delta\pi$ than others. Re-weighting fewer heads fails to fully alleviate the suppression from RAG-suppression heads, while exceeding this optimal number can harm the performance of non-RAG-suppression heads, ultimately diminishing overall effectiveness.

### v2

@@ Abstract
Large language models (LLMs) enhanced with retrieval-augmented generation (RAG) have introduced a new paradigm for web search. However, the limited context awareness of LLMs degrades their performance on RAG tasks. Existing methods to enhance context awareness are often inefficient, incurring time or memory overhead during inference, and many are tailored to specific position embeddings. In this paper, we propose Position-Embedding-Agnostic attention Re-weighting (PEAR), which enhances the context awareness of LLMs with zero inference overhead. Specifically, on a proxy task focused on context copying, we first detect heads which suppress the models’ context awareness, thereby diminishing RAG performance. To weaken the impact of these heads, we re-weight their outputs with learnable coefficients. The LLM (with frozen parameters) is optimized by adjusting these coefficients to minimize loss on the proxy task. During inference, the optimized coefficients are fixed to re-weight these heads, regardless of the specific task at hand. Our proposed PEAR offers two major advantages over previous approaches: (1) It introduces zero additional inference overhead in terms of memory usage or inference time, while outperforming competitive baselines in accuracy and efficiency across various RAG tasks. (2) It is independent of position embedding algorithms, ensuring broader applicability.
@@ 1 Introduction
Retrieval-augmented generation (RAG, (Lewis et al., 2021)) is widely utilized to enhance large language models (LLMs) on tasks like question answering. Typically, an RAG framework retrieves documents related to users’ questions from external knowledge bases or web pages, and then arranges them in the LLMs’ context as the references to form answers. This LLM-based question-answering paradigm has given rise to a promising web search paradigm (Microsoft, 2023; OpenAI, 2024).
Recent research demonstrated LLMs’ limitations on context awareness, especially when processing long context. These limitations in LLMs’ context awareness challenge the effectiveness and robustness of RAG frameworks. For instance, Liu et al. (Liu et al., 2023) found that when performing in-context retrieval tasks, LLMs exhibit insensitivity to information located in the middle of the context, a phenomenon referred to as “lost in the middle.” Chen et al. (Chen et al., 2024) identified a mathematical property in rotary position embedding (RoPE, (Su et al., 2023)) that results in LLMs assigning less attention to specific contextual positions, leading to varying context awareness throughout the entire context.
Existing approaches to enhancing LLMs’ context awareness are inefficient in terms of memory and time cost. Some works (Peysakhovich & Lerer, 2023) segment and re-arrange the input context, with the assumption that placing important information in positions the model attends well can improve RAG’s effectiveness. This method incurs additional inference time costs, negatively affecting user experience, as it requires multiple forward passes to obtain attention weights for guiding segment rearrangement. Another group of studies modifies the model’s working mechanism, specifically by employing a set of position embeddings to adjust the attention preferences of attention heads. While these methods are input-agnostic, they also lack efficiency due to “parallelable” forward passes (Chen et al., 2024) with slightly increased time cost in additional aggregation operation, or disrupt the parallelism of multi-head attention (resulting in increased time cost) to achieve low memory cost (Lin et al., 2024), alternatively, still requiring “non-parallelable” multiple forward passes (Zhang et al., 2024) which is similar to (Peysakhovich & Lerer, 2023). Moreover, these studies are mainly designed for RoPE and face challenges in generalizing to other position embedding algorithms, limiting their broader applications.
@@
    Prior research (McDougall et al., 2023; Lv et al., 2024a) detected some attention heads decreasing the language model’s prediction confidence by suppressing the flow of contextual information to the final position within the context, where the output is to be generated.
@@
As a result, we contend that such suppression mechanism in LLMs can be safely111“Safely” means that the parametric knowledge and fundamental capabilities remain unaffected. Detailed experimental results are presented in Section 5.4. weakened in RAG scenarios, thereby improving the RAG performance of LLMs. Our proposed PEAR includes two stages:
@@
In the second stage, we weaken detected RAG-suppression heads by re-weighting their outputs using learnable coefficients. These coefficients are optimized by minimizing the LLM’s loss on the proxy task, with the objective of next-token prediction in a supervised fine-tuning process (loss is computed only for the second half of the random-token sequence). During the optimization, the original LLM parameters are frozen. Intuitively, most of the learned coefficients are optimized to values less than one, reducing the relative weight of these heads compared to others in the same layer when multi-head outputs are aggregated. Consequently, their influence during the forward pass is weakened. Once optimized, the coefficients remain fixed and are agnostic to downstream RAG tasks.
PEAR achieves two-fold contributions:
@@ 2.1 Context awareness enhancement
Many studies highlighted limitations in LLMs’ context awareness. For example, Lu et al. (Lu et al., 2022) found that the order of in-context learning (ICL) demonstrations significantly affects ICL accuracy. Liu et al. (Liu et al., 2023) demonstrated that LLMs exhibit stronger awareness of content at the beginning and end of context but weaker awareness in the middle, a phenomenon termed “lost in the middle.” Chen et al. (Chen et al., 2024) proposed that LLMs’ context awareness fluctuates across token positions due to mathematical properties in rotary position embedding. These challenges impact applications like RAG that rely on robust context awareness.
@@
Our proposed method, PEAR, enhances context awareness by weakening the RAG-suppression heads during the forward pass. It introduces no additional modules or extra forward passes, resulting in zero additional overhead in memory usage and inference time. Additionally, PEAR operates independently of position embedding algorithms, making it applicable to more LLMs compared to existing approaches.
@@ 2.2 Mechanistic interpretability
Investigating the role of a specific head during a forward pass is one of key focuses in mechanistic interpretability research. Wang et al. (Wang et al., 2023) reported that in GPT-2 small (Radford et al., 2019), the 7th head in the 10th layer, termed the “negative head,” significantly hinders answer copying from context. McDougall et al. (McDougall et al., 2023) comprehensively studied this head, suggesting it functions as a self-repair mechanism to prevent overconfident outputs. Lv et al. (Lv et al., 2024a) found that negative heads exist across various LLMs, employing different mechanisms to mitigate overconfidence, such as generating counteracting vectors or introducing high-frequency tokens’ information. This paper does not examine what specific mechanisms the heads employ to suppress RAG performance but instead aims to discover and suppress heads negatively impactful across general RAG tasks.
Yu et al. (Yu et al., 2023) detected two types of heads in Transformer-based language models during counter-factual task execution (where counter-factual knowledge is provided in the context): memory heads, which prefer to use stored knowledge, and in-context heads, which prefer to use facts in the context. However, when re-weighting these heads, only reducing the weight of memory heads successfully enhances the model’s preference of using contextual knowledge, while enhancing or mitigating in-context heads does not bring much influence. Moreover, existing results only demonstrate effectiveness of these heads in the “country-capital” task (i.e., prompting the model to answer the capital city of the given country). Additionally, there is no evidence suggesting that re-weighting these heads improves the comprehensive context awareness of LLMs. We owe these failures to their head detection method, and the same re-weighting value applied for all the heads of the same type. By contrast, in our proposed PEAR, each individual head is re-weighted by a specific learnable coefficient, which is optimized through a proxy task independent of downstream tasks.
@@ 3 Preliminaries: Discovery of influential attention heads
    We conduct an intervention on a particular head $A^{(l,h)}$ at a specific position $i$ (e.g., the country token position in above examples) in the normal run by substituting its outputs with $\tilde{\textbf{a}}^{(l,h)}_{i}$. The subsequent activations in the computational graph are then recomputed (these recomputed activations are denoted as $\bar{\textbf{a}}$ in the figure). If the language model’s final output matches the intervention’s expectation (e.g., the predicted token changes from ”Paris” to ”London”), the head $A^{(l,h)}$ is considered to have a positive influence on the processing of sequence $X$.
@@ 4.1 Discovery of RAG-suppression heads
For each input sample, we create a sequence of length $n$, denoted as $\{x_{1},\ldots,x_{n}\}$, where each $x_{i}$ is a randomly sampled token from the vocabulary. This sequence is repeated to form an input sample $X=\{x_{1},\ldots,x_{2n}\}$, with $x_{i}=x_{i+n}$ for $i\in[1,n]$. Research has shown that, in semantically meaningless contexts, models tend to check if the last few tokens in the sequence appeared previously and copy the suffix of their last appearance as the output (Olsson et al., 2022; Lv et al., 2024b). We consider an arbitrary LLM to successfully perform the proxy task when, at position $n+i$, the token with the highest output logits is $x_{i+1}$. Table 1 shows an example input.
@@
    The random token composition in $X$ ensures semantically meaningless input, minimizing knowledge bias and thereby ensuring the discovered attention heads to have general RAG-related functions, independent of specific downstream tasks.
@@
    We focus on detecting changes in logits at position $12n-$ 1, where the model is expected to copy the token from position $n$. Consequently, we intervene at $\textbf{a}^{(l,h)}_{2n-1}$ by replacing it with the saved mean vectors.
@@
    $\Delta\pi^{(l,h)}=\frac{\tilde{\pi}^{(l,h)}_{2n-1}[x_{n}]}{\pi_{2n-1}[x_{n}]}-1,$ (1)
    where $\pi_{2n-1}$ represents the final logits at position $12n-$ 1 during the normal run, and $[x_{n}]$ denotes selecting the value of the token $x_{n}$ from the logits. $\tilde{\pi}^{(l,h)}$ indicates the logits after intervention on $A^{(l,h)}$. We contend that a higher value of this metric suggests a stronger suppression effect from $A^{(l,h)}$.
@@
Notably, we do not suggest that the heads we discovered suppressing RAG tasks operate through the same mechanisms. In Section 5.6, we demonstrate that these heads may serve various functions, such as copy suppression (McDougall et al., 2023), incorporating high-frequency token information (Lv et al., 2024a), or influencing the behavior of other heads to indirectly affect outputs (Wang et al., 2023). Analyzing the specific mechanism for each head is not the focus of this paper and is left for future works.
@@ 4.2 Re-weighting coefficient learning
In theory, our approach, PEAR, introduces zero additional overhead during inference on downstream RAG tasks, as it does not incorporate extra computational modules; instead, it only adjusts the aggregation weights of specific heads. Additionally, the learning of re-weighting coefficients is independent of the LLM architecture, thus making our method compatible with various position embedding algorithms.
@@ 5.1 Setup
Difference between learning and inference
We conducted internal loading training on the re-weighting coefficients $\tau^{(l,h)}$ during the training process. However, during inference, we externally re-weighting the output matrix weights of specific heads before loading the model. Therefore, our method is simple to train and has zero inference overhead.
@@ 5.3 Applicability to LLMs using various position embeddings
In this section, we demonstrate the applicability of PEAR to LLMs utilizing different position embeddings. We conduct a multi-document question-answering (MDQA) experiment based on data from (Liu et al., 2023), which leverage a subset of NaturalQuestions-Open  (Lee et al., 2019; Kwiatkowski et al., 2019), consisting of 2,655 queries. Each query is paired with a context consisting of 10 documents with an average of 1,722 tokens. Following (Liu et al., 2023; Chen et al., 2024), we position the gold document (i.e., the document contains the ground truth answer) at various contextual positions to evaluate the robustness of a context-awareness enhancement method. In our experiments, we set the maximum document count to 10 and assess the question-answering accuracy when the gold document is placed as the 1st, 3rd, 5th, 7th, and 10th document, respectively. Since baseline methods are not compatible with the OPT and Baichuan models, we compare PEAR only with the original models. The results are presented in Table 5.
@@ 5.4 PEAR does not diminish knowledge capabilities in LLMs
Previous research (Geva et al., 2023; Lv et al., 2024a) has shown that certain attention heads store or play a crucial role in eliciting parametric knowledge. This raises the question of whether PEAR enhances context awareness of LLMs at the expense of their ability to utilize this parametric knowledge.
To investigate this, we evaluated a PEAR-enhanced Llama2-7B-chat model using the MMLU benchmark (Hendrycks et al., 2021), and the results are presented in Table 6. The performance of the enhanced Llama2-7B-chat and the original Llama2-7B-chat did not show a significant difference. Consequently, we argue that PEAR, through its effective head discovery and re-weighting learning approaches, does not compromise the knowledge capabilities of LLMs.
@@
5.5 Analysis: The effect of $K$
While we have demonstrated the effectiveness of PEAR from various angles, a key point for discussion is the role of $K$, representing the number of heads to re-weight.
Using Llama2-7B-chat as a case study, we vary $K$ and observe its impact on PEAR’s performance. Table 7 presents results on RAG tasks, while Table 8 details analysis for MDQA tasks. The findings indicate that PEAR performs optimally when $K$ matches the inherent threshold of the model (i.e., $K=30$ for Llama2-7B-chat), i.e., the number of heads with a significantly higher $\Delta\pi$ than others.
Re-weighting fewer heads fails to fully alleviate the suppression from RAG-suppression heads, while exceeding this optimal number can harm the performance of non-RAG-suppression heads, ultimately diminishing overall effectiveness.
5.6 Analysis: The Value of $\tau$
Figure 4: The learned coefficients of PEAR (on Llama2-7B-chat and $K=40$).
Using Llama2-7B-chat as an example, we present the learned coefficients of PEAR in Figure 4, with heads ranked by their $\Delta\pi$ scores.
Intuitively, most heads are optimized to values less than one, which reduces their relative weight compared to other heads within the same layer when multi-head outputs are aggregated. Due to BF16 training precision, many $\tau$s are optimized to the same value. However, using FP32 precision for training did not significantly impact the results. Notably, $\tau$s for 9 heads, which have relatively low $\Delta\pi$, are greater than one. We do not attribute this to the precision of the discovery process, as constraining the re-weighting coefficients to be less than one led to suboptimal performance. Thus, a plausible explanation is that RAG suppression is a complex, cooperative effect involving multiple heads, each with distinct working mechanisms, as discussed in Section 4.1.